The events A, B, and C occur with respective probabilities 0.79, 0.45, and exclusive; likewise the events C and B are mutually exclusive. The probability of the event Au(Bnc) (If necessary, consult a list of. fomuias-) 0.04. The events C and A are mutually probability of the event AnBis 0.30.Compute the
Answer #1

Please don't hesitate to give a "thumbs up " for the answer, in case you're satisfied with it.

A U (B and Cbar) = ?

P(B and Cbar) = P(B) [why ? since C and B are mutually exclusive, Cbar is super-set of B. So, an intersection between Cbar and B will be B]

So,

P(A U (B and Cbar))

= P(A U B)

= P(A)+P(B) - P(A and B)

= .79+.45-.30

= .94